OpenStudy (anonymous):

helpppp.... The perimeter of a rectangle is 256 inches. How long are the sides if the length exceeds the width by 16 inches? Mark lets x represent the width of the rectangle and writes the equation 2x + 2(x + 16) = 256. Solving for x, the width of the rectangle is 56 60 68

OpenStudy (anonymous):

perimeter = 2*a + 2*b if a=x, then b=x+16

OpenStudy (anonymous):

i don't know what i am subpose to do....

OpenStudy (anonymous):



OpenStudy (anonymous):

@khofer you have to algebraically solve 2x + 2(x + 16) = 256

OpenStudy (nurali):

The perimeter of a rectangle is 256 inches. find the length if it exceeds the width by 16 inches. We need the formula P = 2l + 2w You are given the perimeter to be 256 inches. We are also told that: width = x length = x + 16 Then: 256 = 2(x + 16) + 2x 256 = 2x + 32 + 2x 256 = 2x + 32 256 - 32 = 4x 224 = 4x 224/4 = x 56 = x

OpenStudy (anonymous):

ok i got it thank u sooooo much...

OpenStudy (nurali):

My Pleasure.